OpenStudy (anonymous):

Help - Give An Equestion For The Line That Contains ( 3,2 ) And (5,-2)

OpenStudy (anonymous):

The slope first which would be done by using the formula y2-y1/x2-x1. or -2-2/5-3= -4/2 =-2

OpenStudy (anonymous):

first find the slope. it is \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{-2-2}{5-3}=\frac{-4}{2}=-2\]

OpenStudy (anonymous):

then apply the formula y-y1=m(x-x1)

OpenStudy (anonymous):

then point slope formula \[y-y_1=m(x-x_1)\] \[y-2=-2(x-3)\] then solve for y \[y-2=-2x+6\] \[y=-2x+8\]

OpenStudy (anonymous):

Satellite beat me to it :)

OpenStudy (anonymous):

find slope m = y2 - y1 / x2 - x1 m = -2 - 2/ 5 - 3 m = -4 / 2 m = -2 find b y = mx + b pick any co-ordinate 2 = -2 (3) + b 2 = -6 + b 2 + 6 = b 8 = b your equation is y = -2x + 8